Two independent random samples resulted in the following. Find the estimate for the standard error for the difference between two means. (Give your answer correct to two decimal places.)

Sample A: nA = 26, sA = 8.8
Sample B: nB = 29, sB = 11.3
Answer #1

Solution,

Given that,

nA = 26

sA = 8.8

nB = 29

sB = 11.3

so,

SE ( xA, xB ) = (sA / nA ) + (sB / nB)

SE ( xA, xB ) = (8.8 / 26) + (11.3 / 29)

SE ( xA, xB ) = 0.338461538 + 0.389655172

SE ( xA, xB ) = 0.85